# Energy based diffusion generator for efficient sampling of Boltzmann distributions

### v2

@@ 3.1 Model architecture
Fig. 1 offers a visual depiction of the decoding and encoding processes. It is important to highlight that latent diffusion models, which are primarily designed to address generative modeling in data-driven scenarios, which have recently gained significant attention [46, 47, 48]. Their main idea to use a pre-trained encoder and decoder to obtain a latent space that both effectively represents the data and facilitates efficient sampling, with the diffusion model learning the distribution of latent variables. Our model, EDG, leverages a similar idea to address energy-based sampling problems. The key differences between EDG and previous latent diffusion models in structure and algorithm are as follows: first, in EDG, the diffusion model itself functions as the encoder, eliminating the need for a separate encoder; second, by using a unified loss function, the decoder is trained jointly with the diffusion model (see Sec. 3.2).
Figure 1: Probabilistic graphs of the decoding and encoding processes, highlighting the trainable parts of the model in gray.
@@
Below, we present the construction details of modules in EDG, which are used in our experiments. In practical applications, more effective neural networks can be designed as needed.
@@ 3.1.1 Boundary condition-guided score function model
Considering that the true score function satisfies the following boundary conditions for $t=0,T$:
@@ 3.1.2 Generalized Hamiltonian dynamics-based decoder
Inspired by generalized Hamiltonian dynamics (GHD) [12, 13], the decoder generates the output $x$ using the following process. First, an initial sample and velocity ($y,v$) are generated according to the latent variable $z_{0}$. Then, $(y,v)$ is iteratively updated as follows:
$v$ ::=:= $v-\frac{\epsilon(l;\phi)}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right),$
$y$ ::=:= $y+\epsilon(l;\phi)\left(v_{k}\odot e^{\epsilon_{0}Q_{y}(v_{k},l;\phi)}+T_{y}(v_{k},l;\phi)\right),$
$v$ ::=:= $v-\frac{\epsilon(l;\phi)}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right).$
@@
$x=y-\exp(\epsilon_{0}\eta(y;\phi))\nabla U(y)+2\exp(\epsilon_{0}\eta(y;\phi))\xi$
@@
where $\xi$ is distributed according to the standard Gaussian distribution. The equation can be interpreted as a finite-step approximation of Brownian dynamics $\mathrm{d}y=-\nabla U(y)\,\mathrm{d}t+\mathrm{d}W_{t}$. In the above procedure, $Q_{v}$, $T_{v}$, $Q_{y}$, $T_{y}$, $\epsilon$, and $\eta$ are all trainable neural networks.
@@
The key advantages of the GHD-based decoder are twofold. First, it effectively leverages the gradient information of the energy function, and our experiments show that it can enhance the sampling performance for multi-modal distributions. Second, by incorporating trainable correction terms and steps into the classical Hamiltonian dynamics, it achieves a good decoder density with only a few iterations. The complete decoding process can be found in C.
3.2 Loss function
To optimize parameters of the decoding and encoding processes, we can minimize the KL divergence between the joint distributions of $(x,z_{[\cdot]})$ provided by the decoding and encoding processes. The divergence is also an upper bound of $D_{KL}(p_{D}(x)||\pi(x))$ like that in (1) due to the data processing inequality. As derived in B, the KL divergence can be expressed as
$D_{KL}\left(p_{D}(x,z_{[\cdot]})||p_{E}(x,z_{[\cdot]})\right)=\mathcal{L}(\theta,\phi)+\log Z,$ (5)
where
$+\int_{0}^{T}\frac{g(t)^{2}}{2}\mathbb{E}_{p_{D}}\Bigg{[}\left\|s\left(z_{t},x,t;\theta\right)\right\|^{2}+2\nabla_{z_{t}}\cdot s\left(z_{t},x,t;\theta\right)$
$\hskip 170.71652pt+\left\|\nabla_{z_{t}}\log p_{D}(z_{t})\right\|^{2}\Bigg{]}\mathrm{d}t.$
By utilizing the importance-sampled integration and Hutchinson estimator, we can obtain an equivalent expression of $\mathcal{L}(\theta,\phi)$, which can be efficiently and unbiasedly estimated by Monte Carlo random sampling from $p_{D}$ without numerical solutions to SDEs:
$+\int_{0}^{T}\lambda(t)\cdot\mathbb{E}_{p(t)p(\epsilon)p_{D}(x,z_{t})}\left[\mathcal{L}_{t}(x,z_{t},\epsilon;\theta)\right]\mathrm{d}t,$
where
$\mathcal{L}_{t}(x,z_{t},\epsilon;\theta)=\left\|s\left(z_{t},x,t;\theta\right)\right\|^{2}+2\frac{\partial\left[\epsilon^{\top}s\left(z_{t},x,t;\theta\right)\right]}{\partial z_{t}}\epsilon+\left\|\nabla_{z_{t}}\log p_{D}(z_{t})\right\|^{2},$
$p(\epsilon)$ is the Rademacher distribution with $\mathbb{E}[\epsilon]=0,\mathrm{cov}(\epsilon)=I$, $p(t)$ is a proposal distribution of $t\in[0,T]$, and the weighting function $\lambda(t)=\frac{g(t)^{2}}{2p(t)}$. All neural networks involved in EDG can then be trained by minimizing the loss function (7) using stochastic gradient descent.
In our experiments, to mitigate the variance inherent in the Monte Carlo approximation of the loss function, we adopted the strategy from [49] to dynamically adjust the time proposal distribution $p_{D}(t)$, based on a historical buffer that stores recent loss values related to $t$ in the r.h.s of (7). For more details, please refer to D.
3.3 Sample reweighting
$\mathbb{E}_{\pi(x)}[O(x)]\approx\frac{1}{N}\sum_{n}O(x^{n}).$
and obtain a consistent estimate of $\mathbb{E}_{\pi(x)}[O(x)]$ as:
$\mathbb{E}_{\pi(x)}[O(x)]\approx\frac{\sum_{n}w(x^{n},z_{0}^{n})O(x^{n})}{\sum_{n}w(x^{n},z_{0}^{n})},$
where the estimation error approaches zero as $N\to\infty$.
In addition, the weight function $w$ can also be used to estimate the normalizing constant $Z$, which is a crucial task in many applications, such as Bayesian model selection in statistics and free energy estimation in statistical physics. Based on (1), we have:
$\log Z\geq-\mathbb{E}_{p_{D}(x,z_{0})}\left[\log w(x,z_{0})\right].$
Here, the lower bound can also be estimated using samples from the decoder, and the tightness of this bound is achieved when $p_{D}(x,z_{0})=p_{E}(x,z_{0})$.
The main difficulty in above computations lies in the intractability of the marginal decoder density $p_{E}(z_{0}|x)$ when calculating the weight function. To overcome this, following the conclusion from Sec. 4.3 in [37], we construct the following probability flow ordinary differential equation (ODE):
$\mathrm{d}z_{t}=\left(f(z_{t},t)-\frac{1}{2}g(t)^{2}s(z_{t},x,t;\theta)\right)\mathrm{d}t,$
with the boundary condition $z_{T}\sim p_{E}(z_{T})$. If $s(z_{t},x,t;\theta)$ accurately approximates the score function $\nabla_{z_{t}}\log p_{E}(z_{t}|x)$ after training, the conditional distribution of $z_{t}|x$ given by the ODE will match that of the encoding process for each $t\in[0,1]$. Consequently, we can employ the neural ODE method [50] to efficiently compute $p_{E}(z_{0}|x)$.
@@ 4 Experiments
We conduct an empirical evaluation of EDG across a diverse range of energy functions. First, we present the results obtained from a set of two-dimensional distributions. Next, we showcase the performance of EDG in the context of Bayesian Logistic regression. Last, we apply EDG to an Ising model. All the experimental details are provided in E. Additionally, we perform an ablation study to verify the effectiveness of each module in EDG. Please refer to F for more information.
@@
    Vanilla Hamiltonian Monte Carlo method [8], denoted as V-HMC.
@@
    L2HMC [13], a GHD-based MCMC method with a trainable proposal distribution model.
@@
    Boltzmann Generator (BG) [18], a VI method where the surrogate distribution is modeled using RealNVP [51].
@@
    Neural Renormalization Group (NeuralRG) [17], a method similar to BG and designed specifically for the Ising model. In this section, NeuralRG is only used for experiments of the Ising model.
@@
    Path Integral Sampler (PIS) [39], a diffusion-based sampling model via numerical simulation of an SDE.
Table 1: The Maximum Mean Discrepancy (MMD) between the samples generated by each generator and the reference samples. Details on the calculation of discrepancy can be found in E.
@@
Figure 2: Density plots for 2D energy function. For the generation of reference samples, please refer to E. We generate $500,000$ samples for each method and plot the histogram.
2D energy function Firstly, we compare our model with the other models on several synthetic 2D energy functions: MoG2(i) (Mixture of two isotropic Gaussians with the same $\sigma^{2}=0.5$ or different variance $\sigma_{1}^{2}=1.5,\sigma_{2}^{2}=0.3$ and centroids separated by distance $10$), MoG6 (Mixture of six isotropic Gaussians with variance $\sigma^{2}=0.1$), MoG9 (Mixture of nine isotropic Gaussians with variance $\sigma^{2}=0.3$), Ring, Ring5 (energy functions can be seen in [12]). We present the histogram of samples for visual inspection in Fig. 2, and Table 1 summarizes the sampling errors. As shown, EDG delivers higher-quality samples compared to the other methods. To elucidate the function of each component in EDG, we compare our model with vanilla VAE in F.
@@
based on the training set $\mathcal{D}_{\mathrm{train}}$, where the prior distribution $p(x)$ is a standard Gaussian distribution. Then, for a given $D$, the conditional distribution $p(L|D,\mathcal{D}_{\mathrm{train}})$ can be approximated as $\sum_{n}p(L|D,x^{n})$. We conduct experiments on three datasets: Australian (AU, 15 covariates), German (GE, 25 covariates), and Heart (HE, 14 covariates) [52], evaluating accuracy rate (ACC) and Area Under the Curve (AUC) on the test subset. Remarkably, as illustrated in Table 2, EDG consistently achieves the highest accuracy and AUC performance.
@@
We extend our analysis to the binary Covertype dataset comprising 581,012 data points and 54 features. The posterior of the classifier parameters follows a hierarchical Bayesian model (see Sec. 5 of [27]), with $x$ denoting the combination of classifier parameters and the hyperparameter in the hierarchical Bayesian model. To enhance computational efficiency, in BG and EDG, $\log\pi(x)$ is unbiasedly approximated during training as
@@
Table 3: Classification accuracy on the test dataset of Coverstype. The reported values represent averages and standard deviations of accuracy over 32 independent experiments.
V-HMC L2HMC BG PIS EDG

### v3

@@ 3.1 Model architecture
Figure 1: Illustration of the decoding and encoding processes in EDG. In the decoding process, the latent variable $z_{0}$ is drawn from a tractable prior $p_{D}(z_{0})$, and a sample $x$ is generated via the decoder distribution $p_{D}(x|z_{0})$ (modeled as Gaussian in this work). Simultaneously, a latent path $z_{[\cdot]}=\{z_{t}\}_{t\in[0,T]}$ is generated by a forward SDE. In the encoding process, $x\sim\pi(x)$ and $z_{T}\sim p_{E}(z_{T})$ are assumed to be independently distributed, and a reverse-time SDE models the conditional distribution of the latent path given $x$ and $z_{T}$, where $s(z_{\tilde{t}},x,\tilde{t};\theta)$ approximates the score function $\nabla_{z_{\tilde{t}}}\log p_{D}(z_{\tilde{t}}|x;\phi)$. Given that training samples of $x$ and $z_{t}$ are generated exclusively by the decoding process, while the encoding process is used solely for loss computation (not for actual sampling), we use solid arrows to indicate probabilistic dependencies in the decoding process and dashed arrows for those in the encoding process. Both the decoder parameters (mean and covariance) and the score model are parameterized by neural networks and jointly optimized via the loss function $\mathcal{L}(\theta,\phi)$.
Fig. 1 provides a schematic overview of the decoding and encoding processes in EDG. The decoding process formulated in Eq. (2) can be readily simulated. In contrast, the encoding process defined by Eq. (4) is not directly simulatable, as no samples from the target distribution are available prior to training. However, as will be discussed in Sec. 3.2, by optimizing the parameters of the encoding process, we can leverage samples from the decoding process and the probabilistic model induced by the encoding process to obtain an accurate variational estimate of the KL divergence between the marginal distribution of $x$ produced by the decoder and the target distribution $\pi(x)$. Based on this estimate, we can jointly train both the decoder and the conditional score network in the encoding process, thereby enabling the generation of high-quality samples that closely match the target distribution. Accordingly, the encoding process in EDG functions as a powerful encoder, replacing the conventional parametric encoder typically used in variational methods (see $p_{E}(z_{0}\mid x;\theta)$ in Eq. (1)).
3.2 Loss function
Based on the architecture of the EDG, we establish the following theorem that provides a theoretical foundation for model training. It shows that the KL divergence between the joint distributions of $(x,z_{[\cdot]})$ induced by the two processes provides an upper bound for $D_{\mathrm{KL}}(p_{D}(x)||\pi(x))$. For a fixed decoder $p_{D}(x|z_{0};\phi)$, minimizing the joint divergence with respect to the score model parameters $\theta$ tightens this upper bound. Moreover, jointly optimizing both $\phi$ and $\theta$ based on the joint divergence can effectively reduce the gap between the generated and target distributions of $x$.
Theorem 1
For the decoding and encoding processes defined by Eq. (2) and Eq. (4), we have
$D_{\mathrm{KL}}\left(p_{D}(x)\,\|\,\pi(x)\right)$ $\leq$ $D_{\mathrm{KL}}\left(p_{D}(x,z_{[\cdot]})\,\|\,p_{E}(x,z_{[\cdot]})\right)$ (6)
$=$ $\mathcal{L}(\theta,\phi)+\int_{0}^{T}\frac{g(t)^{2}}{2}\mathbb{E}_{p_{D}}\left[\|\nabla_{z_{t}}\log p_{D}(z_{t})\|^{2}\right]\mathrm{d}t$
$+\log Z,$
where $\mathcal{L}(\theta,\phi)$ is given by
$+\int_{0}^{T}\frac{g(t)^{2}}{2}\mathbb{E}_{p_{D}}\Bigg[\left\|s\left(z_{t},x,t;\theta\right)\right\|^{2}+2\nabla_{z_{t}}\cdot s\left(z_{t},x,t;\theta\right)\Bigg]\mathrm{d}t.$
Moreover, the equality in Eq. (6) holds if $z_{0}$ is independent of $z_{T}$ in the decoding process, and $s(z,x,t;\theta)\equiv\nabla_{z_{t}}\log p_{D}(z_{t}|x;\phi)$.
Proof. See C. $\square$
Since the last two terms in Eq. (6) are independent of the model parameters $\theta$ and $\phi$, we can adopt $\mathcal{L}(\theta,\phi)$ in Eq. (7) as the training loss. To improve computational efficiency, we estimate the divergence term using the Hutchinson estimator and evaluate the time integral via Monte Carlo sampling. This leads to an equivalent expression of $\mathcal{L}(\theta,\phi)$ that allows efficient and unbiased estimation:
$+\mathbb{E}_{t\sim\mathcal{U}[0,T],\epsilon\sim p(\epsilon),(x,z_{t})\sim p_{D}(x,z_{t})}\left[\mathcal{L}_{t}(x,z_{t},\epsilon;\theta)\right],$
where
$\mathcal{L}_{t}(x,z_{t},\epsilon;\theta)=\frac{Tg(t)^{2}}{2}\left(\left\|s\left(z_{t},x,t;\theta\right)\right\|^{2}+2\frac{\partial\left[\epsilon^{\top}s\left(z_{t},x,t;\theta\right)\right]}{\partial z_{t}}\epsilon\right),$ (9)
$\mathcal{U}[0,T]$ denotes a uniform distribution over $[0,T]$, and $p(\epsilon)\sim\mathrm{Rademacher}^{d}$ denotes a random vector in $\mathbb{R}^{d}$ with i.i.d. Rademacher entries, satisfying $\mathbb{E}[\epsilon]=0$ and $\mathrm{Cov}(\epsilon)=I$. This expression enables the use of stochastic gradient descent for parameter optimization. (See D for the proof of equivalence between Eq. (8) and Eq. (7).) The training procedure is outlined in Algorithm 1. It is worth noting that when the SDE in the decoding process follows a standard formulation used in diffusion models (see Eq. (2)), the conditional distribution $p_{D}(z_{t}|z_{0})$ can be sampled exactly, and numerical integration of SDEs is not required during training.
Algorithm 1 Training procedure
for each minibatch do
  for $i=1,\ldots,\text{batch size}$ do
   Sample $z_{0}^{i}\sim p_{D}(z_{0}^{i})$ and $x^{i}\sim p_{D}(x^{i}|z_{0}^{i};\phi)$;
   Sample $t\sim\mathcal{U}[0,T]$, $\epsilon^{i}\sim\text{Rademacher}^{d}$ and $z_{t}^{i}\sim p_{D}(z_{t}^{i}|z_{0}^{i})$;
   Evaluate $\mathcal{L}_{t}(x^{i},z_{t}^{i},\epsilon^{i};\theta)$ using Eq. (9);
   Compute $L_{i}=\log p_{D}(x^{i}|z_{0}^{i};\phi)+U(x^{i})+\mathcal{L}_{t}(x^{i},z_{t}^{i},\epsilon^{i};\theta)$;
  end for
  Compute minibatch loss: $\hat{\mathcal{L}}(\theta,\phi)=\text{average of }\{L_{1},L_{2},\ldots\}$;
  Update parameters: $\phi,\theta\leftarrow\text{Update}(\nabla_{\phi,\theta}\hat{\mathcal{L}}(\theta,\phi))$;
end for
3.3 Sample reweighting
$\mathbb{E}_{\pi(x)}[O(x)]\approx\frac{1}{N}\sum_{n}O(x^{n}).$
and obtain a consistent estimate of $\mathbb{E}_{\pi(x)}[O(x)]$ as:
$\mathbb{E}_{\pi(x)}[O(x)]\approx\frac{\sum_{n}w(x^{n},z_{0}^{n})O(x^{n})}{\sum_{n}w(x^{n},z_{0}^{n})},$ (11)
where the estimation error approaches zero as $N\to\infty$ [neal2001annealed].
The main difficulty in above computations lies in the intractability of the marginal encoder density $p_{E}(z_{0}|x)$ when calculating the weight function. To overcome this, following the conclusion from Sec. 4.3 in [key-2], we construct the following probability flow ODE:
$\mathrm{d}z_{t}=\left(f(z_{t},t)-\frac{1}{2}g(t)^{2}s(z_{t},x,t;\theta)\right)\mathrm{d}t,$ (12)
with the boundary condition $z_{T}\sim p_{E}(z_{T})$. If $s(z_{t},x,t;\theta)$ accurately approximates the score function $\nabla_{z_{t}}\log p_{D}(z_{t}|x)$ after training, the conditional distribution of $z_{t}|x$ given by the ODE will match that of the encoding process for each $t\in[0,T]$. Consequently, we can employ the neural ODE method [ode] to efficiently compute $p_{E}(z_{0}|x)$ as
$\log p_{E}(z_{0}|x)=\log p_{E}(z_{T})+\int_{0}^{T}\nabla\cdot\left(f(z_{t},t)-\frac{1}{2}g(t)^{2}s(z_{t},x,t;\theta)\right)\mathrm{d}t.$ (13)
See E for details of the probability flow ODE calculation, and the reweighting algorithm is summarized in Algorithm 2.
Algorithm 2 Sampling and reweighting procedure
Input: Initial density: $p_{D}(z_{0})$; Decoder: $p_{D}(\cdot|z_{0};\phi^{*})$; Score: $s\left(\cdot;\theta^{*}\right)$; Observation: $O(\cdot)$.
for $i=1,...,\text{sample\ size}$ do
  Sample $z_{0}^{i}\sim p_{D}(z_{0}^{i})$ and $x^{i}\sim p_{D}(x^{i}|z_{0}^{i};\phi^{*})$;
  Compute $O_{i}=O(x^{i})$;
  Compute $p_{E}(z_{0}^{i}|x^{i};\theta^{*})$ by Eqs. (12, 13);
  Compute $w_{i}=\frac{\exp(-U(x^{i}))p_{E}(z_{0}^{i}|x^{i};\theta^{*})}{p_{D}(z_{0}^{i})p_{D}(x^{i}|z_{0}^{i};\phi^{*})}$;
end for
Estimate $\mathbb{E}_{\pi}[\hat{O}(x)]=\frac{\sum_{i}w_{i}O_{i}}{\sum_{i}w_{i}}$;
Output: $\{x_{i}\}$, $\mathbb{E}_{\pi}[\hat{O}(x)]$
In addition, the weight function $w$ can also be used to estimate the normalizing constant $Z$, which is a crucial task in many applications, such as Bayesian model selection in statistics and free energy estimation in statistical physics. Based on Eq. (1), we have:
$\log Z\geq\mathbb{E}_{p_{D}(x,z_{0})}\left[\log w(x,z_{0})\right].$ (14)
Here, the lower bound can also be estimated using samples from the decoder, and the tightness of this bound is achieved when $p_{D}(x,z_{0})=p_{E}(x,z_{0})$. In practice, the estimate of $\log Z$ given by the above inequality can serve as an indicator of training quality. A larger value suggests a more accurate approximation and, consequently, better model performance. More detailed analysis is provided in F.
3.4 Network design
@@
Below, we present the construction details of modules in EDG used in our experiments. In practical applications, more effective neural networks can be designed as needed.
@@ 3.4.1 Boundary condition-guided score function model
Considering that the true score function satisfies the following boundary conditions for $t=0$ and $T$:
@@ 3.4.2 Generalized Hamiltonian dynamics-based decoder
Inspired by generalized Hamiltonian dynamics (GHD) [mc-4, mc-5], the decoder generates the output $x$ using the following process. First, an initial sample and velocity ($y,v$) are generated according to the latent variable $z_{0}$. Then, $(y,v)$ is iteratively updated as follows:
$v$ $:=$ $v-\frac{\epsilon_{0}e^{\epsilon_{0}\epsilon(l;\phi)}}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right),$
$y$ $:=$ $y+\epsilon_{0}e^{\epsilon_{0}\epsilon(l;\phi)}\left(v_{k}\odot e^{\epsilon_{0}Q_{y}(v_{k},l;\phi)}+T_{y}(v_{k},l;\phi)\right),$
$v$ $:=$ $v-\frac{\epsilon_{0}e^{\epsilon_{0}\epsilon(l;\phi)}}{2}\left(\nabla U(y)\odot e^{\frac{\epsilon_{0}}{2}Q_{v}(y,\nabla U(y),l;\phi)}+T_{v}(y,\nabla U(y),l;\phi)\right).$
@@
$x=y-\epsilon_{0}e^{\epsilon_{0}\eta(y;\phi)}\nabla U(y)+\sqrt{2\epsilon_{0}e^{\epsilon_{0}\eta(y;\phi)}}\xi,$
@@
where $\xi$ is distributed according to the standard Gaussian distribution. The equation can be interpreted as a finite-step approximation of Brownian dynamics $\mathrm{d}y=-\nabla U(y)\,\mathrm{d}t+\mathrm{d}W_{t}$. The decoder density is then defined as
$p_{D}(x|z_{0};\phi)=\mathcal{N}\left(x\mid y-\epsilon_{0}e^{\epsilon_{0}\eta(y;\phi)}\nabla U(y),\,2\epsilon_{0}e^{\epsilon_{0}\eta(y;\phi)}I\right),$
where $I$ denotes the identity matrix. Since this density can be tractably evaluated given $x$, $z_{0}$, and $\phi$, all decoder parameters $\phi$ involved in the GHD module can be jointly optimized with the encoder parameters $\theta$ by minimizing the loss function $\mathcal{L}(\theta,\phi)$. In the above procedure, $Q_{v}$, $T_{v}$, $Q_{y}$, $T_{y}$ are all trainable neural networks. To mitigate the impact of step sizes on model performance, we parameterize them as trainable positive functions of the form $\epsilon_{0}\exp(\epsilon_{0}\epsilon(l;\phi))$ and $\epsilon_{0}\exp(\epsilon_{0}\eta(y;\phi))$, where $\epsilon_{0}$ is a small positive constant that controls the initial scale and prevents instability during training caused by large step sizes. The choice of $\epsilon_{0}$ in practice is discussed in J.
@@
The key advantages of the GHD-based decoder are twofold. First, it effectively leverages the gradient information of the energy function, and our experiments show that it can enhance the sampling performance for multi-modal distributions. Second, by incorporating trainable correction terms and steps into the classical Hamiltonian dynamics, it achieves a good decoder density with only a few iterations. The complete configurations can be found in G.
@@ 4 Experiments
We conduct an empirical evaluation of EDG across a diverse range of energy functions. We begin with experiments on a set of two-dimensional distributions, followed by results on Bayesian logistic regression. We then assess EDG’s performance on two Lennard-Jones tasks, and finally apply it to the Ising model. All the experimental details are provided in H. Additionally, we perform an ablation study to verify the effectiveness of each module in EDG. Please refer to J for more information.
@@
    Vanilla Hamiltonian Monte Carlo method [mc-2], denoted as V-HMC.
@@
    L2HMC [mc-5], a GHD-based MCMC method with a trainable proposal distribution model.
@@
    Boltzmann Generator (BG) [key-4], a VI method that generates samples solely based on a prescribed energy function, without relying on any real data. The surrogate distribution is modeled using RealNVP [flow].
@@
    Neural Renormalization Group (NeuralRG) [ising], a method similar to BG and designed specifically for the Ising model. In this section, NeuralRG is only used for experiments of the Ising model.
@@
    Path Integral Sampler (PIS) [diff-6], a diffusion-based sampling model via numerical simulation of an SDE.
We now present the experimental results for each sampling task in detail. An analysis of the computational cost is provided in I.
Table 1: The Maximum Mean Discrepancy (MMD) between the samples generated by each generator and the reference samples. Details on the calculation of discrepancy can be found in H.
@@
2D energy function Firstly, we compare our model with the other models on several synthetic 2D energy functions: MoG2 (Mixture of two isotropic Gaussians with equal variance $\sigma^{2}=0.5$, centers $10$ apart), MoG2(i) (Mixture of two isotropic Gaussians with unequal variance $\sigma_{1}^{2}=1.5,\sigma_{2}^{2}=0.3$, centers $10$ apart), MoG6 (Mixture of six isotropic Gaussians with variance $\sigma^{2}=0.1$), MoG9 (Mixture of nine isotropic Gaussians with variance $\sigma^{2}=0.3$), Ring, Ring5 (energy functions can be seen in [mc-4]). We present the histogram of samples for visual inspection in Fig. 2, and Table 1 summarizes the sampling errors. As shown, EDG delivers higher-quality samples compared to the other methods. To further clarify the contribution of each component in EDG, we compare it with a vanilla VAE in J. The effectiveness of the reweighting scheme is also evaluated in H.
@@
Figure 2: Density plots for 2D energy function. For the generation of reference samples, please refer to H. We generate $500,000$ samples for each method and plot the histogram.
@@
based on the training set $\mathcal{D}_{\mathrm{train}}$, where the prior distribution $p(x)$ is a standard Gaussian distribution. Then, for a given $D$, the conditional distribution $p(L|D,\mathcal{D}_{\mathrm{train}})$ can be approximated as $\frac{1}{N}\sum_{n}p(L|D,x^{n})$. We conduct experiments on three datasets: Australian (AU, 15 covariates), German (GE, 25 covariates), and Heart (HE, 14 covariates) [logistic1], evaluating accuracy rate (ACC) and Area Under the Curve (AUC) on the test subset. Remarkably, as illustrated in Table 2, EDG consistently achieves the highest accuracy and AUC performance.
@@
Table 3: Classification accuracy on the test dataset of Covertype. The reported values represent averages and standard deviations of accuracy over 32 independent experiments.
V-HMC L2HMC BG PIS EDG
We extend our analysis to the binary Covertype dataset comprising 581,012 data points and 54 features. The posterior of the classifier parameters follows a hierarchical Bayesian model (see Sec. 5 of [vi-7]), with $x$ denoting the combination of classifier parameters and the hyperparameter in the hierarchical Bayesian model. To enhance computational efficiency, in BG and EDG, $\log\pi(x)$ is unbiasedly approximated during training as
@@
Lennard-Jones We further evaluate the performance of EDG on two Lennard–Jones (LJ) systems with non-periodic boundary conditions, consisting of 13 and 55 particles (LJ13 and LJ55), as described in [klein2023equivariant]. The LJ potential is a commonly used interaction model that captures both repulsive and attractive forces between non-bonded particles. The potential energy depends on the pairwise distances between particles, and its explicit form is provided in H. In this task, we temporarily disregard the issue of model equivariance, which we leave for future investigation.
LJ13 refers to the system of 13 particles, $x=\{x_{1},...,x_{13}\}$ with 3 dimensions each, resulting in a task with dimensionality $d=39$. LJ55 meanwhile refers to a system of 55 particles, $x=\{x_{1},...,x_{55}\}$ with 3 dimensions each, resulting in a high dimensional task with dimensionality $d=165$. For the experimental results, we evaluate the generated sample with the test data from [klein2023equivariant] obtained by MCMC. A visualization of the interatomic distances is presented in Fig. 3. In H, we test the effectiveness of sample reweighting introduced in Sec. 3.3, by using the relative Effective Sample Size (see Eq. (17) and Table 6).
Figure 3: Comparison of the test data interatomic distance of LJ13 (left), LJ55 (right) with samples generated from BG and EDG. The results for L2HMC and PIS are provided separately in H due to significant deviations from the test data.